OpenStudy (anonymous):

Use the distributive property 4/7(1/9x+35)

OpenStudy (zehanz):

Do you know what the distributive property is?

OpenStudy (anonymous):

Not really..

OpenStudy (zehanz):

a(b+c)=ab+ac

OpenStudy (zehanz):

So:\[\frac{ 4 }{ 7 }\left( \frac{ 1 }{ 9 }x+35 \right)=\frac{ 4 }{ 7 } \cdot \frac{ 1 }{ 9 }x+\frac{ 4 }{ 7 }\cdot 35= ...\]

OpenStudy (anonymous):

Thank you

OpenStudy (zehanz):

YW!